OpenStudy (anonymous):

The radius of a spherical watermelon is growing at a constant rate of 2 centimeters per week. The thickness of the rind is always one tenth of the radius. The volume of the rind is growing at the rate _________________ cubic centimeters per week at the end of the fifth week. Assume that the radius is initially zero.

OpenStudy (phi):

first, what is r after 5 weeks ?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

since the rate is 2. therefore you have to do 2*5

OpenStudy (phi):

yes. next use V = 4pi/3 r^3 and find dV/dt as a function of r and dr/dt

OpenStudy (anonymous):

dV/dt = 4pir^2(dr/dt) i'm not sure if that's right

OpenStudy (phi):

yes, that is correct. dr/dt is given as a constant 2 cm/week so \[ \dot{V} = 4 \pi r^2 \cdot 2\] now we need a value for r (of the "flesh" part, which excludes the rind) the rind is 1/10 of r, which means the flesh has a radius of 0.9r use that : \[ \dot{V} =8 \pi (0.9 r)^2 8\] where r=10 is the entire radius

OpenStudy (anonymous):

i get the first formula but not the second. how did you get the second 8?

OpenStudy (anonymous):

ok so when you differentiate V˙=4πr^2⋅2 you get 16pir right? so how did you get your formula?

OpenStudy (phi):

sorry, the second 8 is a typo

OpenStudy (phi):

\[ \dot{V} =8 \pi (0.9 r)^2 \]

OpenStudy (anonymous):

oh ok so what do i have to do next again?

OpenStudy (phi):

using your formula dV/dt = 4pir^2(dr/dt) with dr/dt =2 r = 0.9r (to adjust because we want to exclude the rind) and r=10 we get dV/dt = 4*pi * (0.9*10)^2 * 2 now multiply

OpenStudy (anonymous):

i got 2035.75204

OpenStudy (phi):

that should be the answer in cubic cm per week

OpenStudy (anonymous):

oh ok. idk if my friend's answer's right, but she got 681.09 cm^3. i also googled the question and the person 680.75 cm^3/week but then again i don't know if they're right

OpenStudy (phi):

do you have a link to the 680.75 answer ?

OpenStudy (phi):

ok, I was doing the volume of the flesh. so we should redo the problem V of the rind is the total volume - volume of the inside (flesh) \[ V= \frac{4}{3} \pi r^3 - \frac{4}{3} \pi (0.9r)^3 \] differentiate that

OpenStudy (phi):

we can first simplify that a bit: \[ V = \frac{4\pi}{3}(1-0.729) r^3 \\ V = \frac{4\pi}{3}\cdot 0.271\ r^3 \]

OpenStudy (anonymous):

how did you get this part? (1−0.729)

OpenStudy (phi):

rewrite (0.9r)^3 as 0.9^3 r^3 = 0.729 r^3 then factor out 4pi/3 r^3

OpenStudy (phi):

you should get 681.097 which rounds to 681.10

OpenStudy (anonymous):

ok i got that. thank you!

OpenStudy (phi):

sorry for the long detour!

OpenStudy (anonymous):

no it's okkk. don't worry :)

OpenStudy (anonymous):

Refer to the attached Mathematica 9 calculation.

OpenStudy (phi):

@robtobey Your derivative of V is with respect to r you must adjust it by multiplying by dr/dt (apply the chain rule) to find dV/dt (i.e. volume change per unit time (a week in this case) )